OpenStudy (anonymous):

4x2 + 9y2 = 72 x - y2 = -1 Select all of the following that are solutions to the system shown. (3, 2) (3, -2) (-3, 2) (-3, -2)

OpenStudy (austinl):

Ok, I would use substitution for this one. Do you know how to start that?

OpenStudy (jdoe0001):

so you have the "x" and the "y" terms squared if they're both POSITIVE, is an ellipse, or a circle which is really just a "round ellipse" if they're both NEGATIVE, is a hyperbola they're both POSITIVE, thus is an ellipse

OpenStudy (jdoe0001):

well, I take that back, IF ONLY ONE OF THEM IS NEGATIVE, then is a hyperbola :)

OpenStudy (anonymous):

i got (3, 2) and (3, -2)...is that right??

OpenStudy (jdoe0001):

ohh, wait.. is a system of equations, my bad heheh

OpenStudy (jdoe0001):

thought they were asking for the center pfft

OpenStudy (anonymous):

x - y2 = -1 -y2=-x - 1 y^2= x+1

OpenStudy (anonymous):

(3, 2) 4=4 good (3, -2) 4=4 good (-3, 2) 4=-2 no (-3, -2) 4=-2 no a and b

OpenStudy (anonymous):

thank you (: